(3.3.11) Consider a value to be significantly low if its z score less than or equal to -2 or consider a value to be significantly high if its z score is greater than or equal to 2.
A data set lists weights​ (grams) of a type of coin. Those weights have a mean of 5.28353 g and a standard deviation of 0.06149 g. Identify the weights that are significantly low or significantly high.
What weights are significantly​ low? Select the correct answer below and fill in the answer​ box(es) to complete your choice.
A. Weights that are less than?
B. Weights that are between? and?
C. Weights that are greater than?
1
Expert's answer
2020-11-22T18:06:46-0500

"\mu=5.28353"

"\sigma = 0.06149"


"Z=\frac{x-\mu}{\sigma}"


"Z \le -2" significantly low values:


"-2=\frac{x-5.28353}{0.06149} \implies x=5.16055"



"Z \geq 2" significantly high values:


"2=\frac{x-5.28353}{0.06149} \implies x=5.40651"


Answer:

weights less than 5.16055 are significantly​ low,

weights more than 5.40651 are significantly​ high.